OpenStudy (anonymous):

Solve this system of linear equations. Separate the x- and y-values with a comma. I.E. (x,y) 14x = -22 - 5y 5x = 60 + 5y @kelliegirl33 @ash2326

OpenStudy (anonymous):

14x = -22 - 5y 5x = 60 + 5y --------------add 19x = 38 x = 19 5x = 60 + 5y 5(19) = 60 + 5y 95 = 60 + 5y 95 - 60 = 5y 35 = 5y 7 = y check.. 5x = 60 + 5y 5(19) = 60 + 5(7) 95 = 60 + 35 95 = 95 (correct) x = 19 and y = 7

OpenStudy (anonymous):

So the correct answer is (19,7)?

OpenStudy (anonymous):

yes...the answer is (19,7) thank you skiller....for the testimonial :)

OpenStudy (anonymous):

You're welcome, and for some reason, it states "Incorrect". Would you be able to know why? >.<

OpenStudy (anonymous):

let me try it a different way...hold on

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (anonymous):

14x = -22 - 5y change to 14x + 5y = -22 5x = 60 + 5y change to 5x - 5y = 60 14x + 5y = -22 5x - 5y = 60 ------------- 19x = 38 x = 2......thats where I messed up the first time.....I put 19 instead of 2 5x - 5y = 60 5(2) - 5y = 60 10 - 5y = 60 -5y = 60 - 10 -5y = 50 y = -10 check... 5x - 5y = 60 5(2) - 5(-10) = 60 10 + 50 = 60 60 = 60 (correct) my bad...a stupid mistake.....answer is (2,-10)

OpenStudy (anonymous):

Okay, thanks :)

OpenStudy (anonymous):

your welcome :)